OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (5, 5) m = -3

OpenStudy (anonymous):

use point slope form.... (y-y1)=m(x-x1) where m is the slope, and (x1,y1) is a point... then Slove for y to get it into y=mx+b form. Want me to evaluate more? or do you got it from here?

OpenStudy (anonymous):

can you evaluate more please?